OpenStudy (anonymous):

In the below system, solve for y in the first equation. x + 3y = 6 2x – y = 10 Thanxx =]

hero (hero):

x + 3y = 6 3y = 6 - x y = (6 - x) /3 I hope that's what they're asking for.

OpenStudy (anonymous):

Yes , Thanxx =]